The mean and standard deviation of a random sample of n measurements are equal to 34.2 and 3.7 respectively.

a. Find a 90% confidence interval for mu if n = 144

( __ , ___) round to three decimal places as needed

b. Find a 90% confidence interval for mu if n = 576

(___, ___) round to three decimal places as needed

c. Find the widths of the condfidence intervals found in parts a and b. What is the effect on the width of a confidence interval of quadrupling the sample size while hold the confidence coefficient fixed?

Option a Quadrupling the sample size while holding the confidence coefficient fixed decreases the width of the confidence interval by a factor of 4.

option b Quadrupling the sample size while holding the confidence coefficient fixed does not affect the width of the confidence interval.

option c Quadrupling the sample size while holding the confidence coefficient fixed decreases the width of the confidence interval by a factor of 2.

option d Quadrupling the sample size while holding the confidence coefficient fixed increases the width of the confidence interval by a factor of 4.

option e Quadrupling the sample size while holding the confidence coefficient fixed increases the width of the confidence interval by a factor of 2.
Answer #1

a)

Standard error of the mean = SEM = S/?N = 0.308

t(?, N-1) = 1.656

Confidence interval = m +/- (t(?, N-1)*SEM)

Mean = 34.2

Lower bound: 33.69

Upper bound: 34.71

b)

Standard error of the mean = SEM = S/?N = 0.154

t(?, N-1) = 1.648

Confidence interval = m +/- (t(?, N-1)*SEM)

Mean = 34.2

Lower bound: 33.95

Upper bound: 34.45

c.

Width in part a. = 34.71-33.69 = 1.02

Width in part b = 34.45-33.95 = 0.5

option c Quadrupling the sample size while holding the confidence coefficient fixed decreases the width of the confidence interval by a factor of 2.